OpenStudy (anonymous):

Complete the table and choose the solution for each give vaule of x in questions 8-13. y=-x2+2

OpenStudy (anonymous):

When x=-2 a: y=-(-2)^2+2=-2; (-2, -2) b: y=(-2)^2+2=2; (-2, 2) c: y=-(2)^2+2=-2; (2, -2) d: y=-(-2)^2-2=2; (-2, 2)

OpenStudy (anonymous):

I just asked the exact same questuib you did! I an soo lost! :P

OpenStudy (anonymous):

*am*

OpenStudy (anonymous):

*question* not doing so great with typing today haha